A shunting locomotive provides impulse of 40kN to set motion a stationary 8 tones wagon which then moves off freely at velocity u against a track resistance of 60N/tone and finally reaches a velocity v after 20 seconds. Find the values of u and v.
1
Expert's answer
2020-06-01T13:10:18-0400

"p=40 kN\cdot s"


"p=m\cdot u \to" "u=\frac{p}{m}=\frac{40000}{8000}=5m/s"


"F=ma\to a=F/m=(60\cdot 8)/8000=0.06m/s^2"


"v=u-at=5-0.06\cdot 20=3.8m/s"


Answer. "u=5" "m\/s" ; "v=3.8" "m\/s"